# Supernematic

### v1

@@ VI.2 Finite temperature
$S_{0}$ is governed by the effective action obtained by taking the classical limit of our quantum effective theory (ignoring the $\tau$-dependent terms in Eqs. 21 and 22),
@@
Here we only consider monomer hopping terms mapped to the boson hopping between nearest neighbors, $tb_{i}^{\dagger}b_{j}$. The higher-order hopping terms can be written in a similar way. $m_{\alpha,i}$ denotes the annihilation operator for a “bionic” monomer residing at plaquette center $i$, and $\alpha,\delta,\gamma=r,g,b$ label the color indices (see Fig. 7 d)(i)) with the following rule. A monomer $m_{g,i}$ located on a green plaquette violates the cluster constraint on that plaquette and, consequently, the dimer constraints on both the red and blue bonds in the corresponding dimer representation, similarly for $m_{r,i}$ and $m_{b,i}$. Interpreting the dimer constraints as Gauss laws, $m_{g,i}$ carries opposite gauge charges under the emergent fields $A^{r}$ and $A^{b}$ (defined on different links as shown in Fig. 7 d)(i)), realizing the “bionic” character introduced in Ref. [13]. The magnetic flux defined by summing the gauge fields $A_{ij}^{\alpha}$ around a hexagonal plaquette centered at site $l$ is related to the conjugate height field as $\sum_{\langle ij\rangle\in\hexagon_{l}}A_{ij}^{\alpha}=3\,\theta_{l}^{\alpha}$.
@@
Therefore, the ordering of the phase angle $\theta_{0}$ is discrete and long-ranged, characterized by a $Z_{3}$ order. The connected two-point correlator exhibits a finite correlation length, distinguishing the thermal phase of the SN from a genuine superfluid phase of $\tilde{N}$ in the thermodynamic limit. Nevertheless, the effect of the threefold anisotropy can remain weak over a large length scale, owing to the small Boltzmann factor $e^{-3\beta U}$ and the high-order hopping amplitude $w_{3}$. The length scale beyond which the anisotropy becomes relevant defines a pinning length $l_{p}$. Within the range of $l_{p}$, the thermal SN phase effectively behaves as a superfluid of $\tilde{N}$. To leading order in $w_{3}$, the pinning length is estimated from the renormalization group analysis[30] as,
@@
Now let us discuss the thermal melting of the SN. Because both $\theta_{0}$ and $\vec{\eta}$ order in SN, it can melt in two ways (Fig.8). A two-stage melting proceeds via an intermediate phase whose nature depends on which order melts first: a) if $\theta_{0}$ disorders while $\vec{\eta}$ remains ordered, the intermediate phase is nematic that breaks $\mathcal{C}_{3}$ but preserves translation; b) if $\vec{\eta}$ disorders first but $\theta_{0}$ retains $Z_{3}$ phase coherence, we call the intermediate phase “super”-fluid due to the phase coherence. Since $\cos(3\theta_{0})$ is relevant at the would-be BKT transition of a true $\tilde{N}$ superfluid [31], we expect that thermal transition of $\theta_{0}$ to be three-state Potts-like. Therefore, in either route, both transitions, between SN and the intermediate phase, and between the intermediate and the fully disordered phase, are in the three-state Potts universality class, reflecting $Z_{3}$ symmetry breaking at each step.
@@ VII Other fillings
Extending beyond $\nu=2/3$, the cluster constraint $\sum_{i\in\varhexagon_{r}}n_{i}\in\{0,1,2,3,4,5,6\}$ yields the “commensurate” fillings $\nu=0,1/3,2/3,1,4/3,5/3,2$. Since $\nu$ and $2-\nu$ are related by particle–hole symmetry and $\nu=0$ is trivial, it suffices to analyze $\nu=1/3$, $2/3$ and $1$. At $\nu=1/3$, Thm.  1 applies directly (no need for the tile set $\Sigma_{2}$), so $\tilde{N}$ remains an exact symmetry of the constrained Hilbert space. In contrast, at $\nu=1$ the leading ring exchange breaks $\tilde{N}$ (Fig.  9a). A constructive scheme (see SM Sec. C) can be proofed to generate all local operators order by order, yielding the leading ring exchanges (Fig.  9b) to be $\mathcal{O}_{1}$ for $\nu=1$, $\mathcal{O}_{2}$ for $\nu=2/3$, and $\mathcal{L}_{2}$ (the “Lemniscate operator” in Ref.[12]) for $\nu=1/3$. Notably, $\mathcal{L}_{2}$ conserves a sublattice dipole $\tilde{\vec{P}}=\sum_{i}n_{A}(x_{i})\vec{x}_{i}-\sum_{j}n_{A}(x_{j})\vec{x}_{j}$, which is violated by higher-order terms such as $\mathcal{L}^{\prime}$. Therefore we also expect the SN phase to exist at $\nu=1/3$. For $\nu=1$, the $\mathcal{O}_{1}$ term gaps the system into a sublattice charge-ordered state.
@@ VIII Summary and discussion
In summary, we uncovered a robust mechanism by which geometric frustration and local Hilbert-space constraints enforce an emergent, globally conserved quantum number for bosons under cluster charging interaction on the honeycomb lattice. By mapping the constrained Hilbert space at $\frac{2}{3}$ ($\frac{4}{3}$) or $\frac{1}{3}$ ($\frac{5}{3}$) filling to a nontrivial tiling problem and invoking the Conway–Lagarias boundary invariant, we rigorously established global conservation of the sublattice polarization $\tilde{N}=N_{A}-N_{B}$, beyond the previously shown $U(1)\times U(1)$ gauge symmetry of classical ground state[13]. By mapping the local constraints to coupled double-dimers, we derive an effective field theory that faithfully incorporates the conservation of $\tilde{N}$ and the local constraints of the bosonic quantum Hamiltonian. Quantum fluctuations within this manifold give rise to an insulating yet macroscopically phase-coherent and gapless “supernematic” (SN) phase, characterized by a unique intertwined order that spontaneously breaks both the emergent global $U(1)$ symmetry and the lattice $C_{3}$ rotation, but not translation. We derived a non-local order parameter representing the SN that can be detected through ODLOR: the “bumped loop” operator, which is markedly different from familiar Wildon loop through its shape-dependent charge under $\tilde{N}$. Furthermore, we showed that the quantum phase transition between the columnar-ordered Mott insulator and the SN phase belongs to the 3D $XY$ universality class. We also analyzed the finite temperature transitions that sequentially restore the broken symmetries.
@@
While explicit derivation of the critical value for the microscopic parameters of the original Hamiltonian is beyond the reach of the effective field theory, a concrete microscopic model for realizing the SN phase will be a trimer variant of the so-called Rokhsar–Kivelson Hamiltonian, dubbed the trimer–Rokhsar–Kivelson (trimer-RK) Hamiltonian. This model is a natural generalization of the RK Hamiltonian introduced for quantum dimer models [5]. The trimer-RK Hamiltonian includes both a ring-exchange term, $\mathcal{O}_{2}$, and a potential term (see Appendix E for the explicit form). At the RK point, where the strengths of these two terms are equal, the ground state is an equal-weight superposition of all trimer-covering configurations, analogous to the dimer model at its RK point. In dimer models on bipartite lattices, such as the honeycomb lattice [35], perturbations away from the RK point toward stronger exchange terms drive the system into an ordered phase. In contrast, quantum fluctuations around the trimer-RK point generate an effective action that stabilizes the SN phase (see Appendix E). The key distinction between the proposed trimer-RK and dimer-RK Hamiltonians is rooted in the $\tilde{N}$ invariance. In the trimer case, the dynamics including stochastic fluctuations must preserve $\tilde{N}$ exactly, leading to a Langevin equation of the same form as that describing volume-conserving surface growth [36]. Upon quantization, this yields the phase-angle Lagrangian identical to Eq. 22, in contrast to the quantum Lifshitz model that governs the dimer RK point on bipartite lattices [37].
@@
As a physical platform to persue the realization of SN, we propose the Rydberg atom array. The Rydberg blockade mechanism has been shown to provide cluster-charging constraints on honeycomb lattices [38]. The algebraic decay of various operators of the SN phase can be directly measured in this platform. Measuring the non-local order parameter can be challenging, but it may be achievable through strategies analogous to that used for topological phases in Ref. [39]. Since the total number of particles is not fixed in Rydberg systems, we need to consider the stability of the SN under doping away from the commensurate fillings ($2/3$ or $1/3$). As was discussed in the thermal violation of $\tilde{N}$ conservation, the monomers induce a threefold anisotropy, which pins the macroscopic phase angle of the SN. Doping also introduces monomers that carry fractionalized charges of $e/3$ [12, 13]. However, since the $\vec{\eta}$ field is pinned in the SN phase, we expect these monomers to be confined at low temperature and to become deconfined only in the “super”-fluid regime at finite temperature.
@@
Acknowledgement. – DM thanks Jiaxin Qiao, Yizhi You, Cenke Xu, Yi-Zhuang You, Jing-Yuan Chen, Zheng-Cheng Gu, T. Senthil for the stimulating discussions. We thank Christopher Mudry and Eduardo Fradkin for feedbacks on the draft. DM is supported by SNSF Grant No. 219339. E-AK was supported in part by the NSF through the grant OAC-2118310 supported by the U.S. Department of Energy through Award Number: DE-SC0023905. This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF10436 to E-AK and grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@ A.2 Symmetry properties under lattice transformations
In the continuum, the coarse-graining introduces non-zero average values of $h^{r,g,b}$, and the lattice symmetry can act with some additional shift on top of the microscopic translation [26]. To see this clearly, we consider the nine colomnar ordering states (flat configurations with constant average colored height fields). The lattice transformation among these COs can be implemented via applying local operators across the lattice, allowing us to derive the values of height fields starting from one specific CO (see Fig.11). Hence, under lattice symmetries, the average heights transform as such,
@@ A.3 Relating coarse-grained fields to microscopic fields
However, this is not the full story due to the smearing in the coarse-graining. To the leading order in vertex operator, this amounts to match the lattice configuration of the colomnar ordering and the corresponding uniform heights in the continuum, analogous to dimer model [26], which yields,
@@ Appendix B Sublattice density structure factor
Now let us consider the structure factor of the sublattice density $\tilde{n}(q)\equiv\sum_{i}n_{i,A}e^{iq\cdot x_{i,A}}-\sum_{j}n_{j,B}e^{iq\cdot x_{j,B}}$. From Eq.39, 41 and 44, we can write $\tilde{n}(q)$ in terms of $h_{0}$ fields. It is readily seen that
@@ Appendix E Perturbing around trimer-RK point
In the quantum dimer model on bipartite lattice, the dimer constraints can be viewed as Gauss law on the sites of the lattices [26], and the low-energy physics can be described by compact $U(1)$ gauge theory in $2+1D$. The quantum fluctutation of the local moves in the dimer model corresponds to the instanton events and leads to confinement. Therefore, the gapless RK point is in general unstable under quantum fluctuation. We can ask the analogous question for the trimer model, that is, starting from the trimer RK point where the wavefunction is an equal weight superposition of all possible classical trimer configuration, what is the role of quantum fluctuation?
@@
The coefficient $K_{\eta}$ can be obtained by matching the Gaussian action of the classical trimer model in Ref. [13]. In the dimer model near RK point, it was realized by Henley [42, 37] that in the long wavelength limit, the plaquette flip dynamics can be captured by a Langevin equation of the height field (similar technique is also under the name of “stochastic quantization”, see [43, 44] and the reference therein). Let us only focus on the $h_{0}$ field. The dynamics of $h_{0}$ follows the following Langevin equation,
@@
which is of the same form as in volume conserving surface growth dynamics[36].

### v2

@@ VI.2 Finite temperature
$S_{0}$ is governed by the effective action obtained by taking the classical limit of our quantum effective theory (ignoring the $\tau$-dependent terms in Eqs. 21 and 23),
@@
Here we only consider monomer hopping terms mapped to the boson hopping between nearest neighbors, $tb_{i}^{\dagger}b_{j}$. The higher-order hopping terms can be written in a similar way. $m_{\alpha,i}$ denotes the annihilation operator for a “bionic” monomer residing at plaquette center $i$, and $\alpha,\delta,\gamma=r,g,b$ label the color indices (see Fig. 7 d)(i)) with the following rule. A monomer $m_{g,i}$ located on a green plaquette violates the cluster constraint on that plaquette and, consequently, the dimer constraints on both the red and blue bonds in the corresponding dimer representation, similarly for $m_{r,i}$ and $m_{b,i}$. Interpreting the dimer constraints as Gauss laws, $m_{g,i}$ carries opposite gauge charges under the emergent fields $A^{r}$ and $A^{b}$ (defined on different links as shown in Fig. 7 d)(i)), realizing the “bionic” character introduced in Ref. [zhang2024bionicfractionalizationtrimermodel]. The magnetic flux defined by summing the gauge fields $A_{ij}^{\alpha}$ around a hexagonal plaquette centered at site $l$ is related to the conjugate height field as $\sum_{\langle ij\rangle\in\hexagon_{l}}A_{ij}^{\alpha}=3\,\theta_{l}^{\alpha}$.
@@
Therefore, the ordering of the phase angle $\theta_{0}$ is discrete and long-ranged, characterized by a $Z_{3}$ order. The connected two-point correlator exhibits a finite correlation length, distinguishing the thermal phase of the SN from a genuine superfluid phase of $\tilde{N}$ in the thermodynamic limit. Nevertheless, the effect of the threefold anisotropy can remain weak over a large length scale, owing to the small Boltzmann factor $e^{-3\beta U}$ and the high-order hopping amplitude $w_{3}$. The length scale beyond which the anisotropy becomes relevant defines a pinning length $l_{p}$. Within the range of $l_{p}$, the thermal SN phase effectively behaves as a superfluid of $\tilde{N}$. To leading order in $w_{3}$, the pinning length is estimated from the renormalization group analysis[kosterlitz1973ordering] as,
@@
Now let us discuss the thermal melting of the SN. Because both $\theta_{0}$ and $\vec{\eta}$ order in SN, it can melt in two ways (Fig.8). A two-stage melting proceeds via an intermediate phase whose nature depends on which order melts first: a) if $\theta_{0}$ disorders while $\vec{\eta}$ remains ordered, the intermediate phase is nematic that breaks $\mathcal{C}_{3}$ but preserves translation; b) if $\vec{\eta}$ disorders first but $\theta_{0}$ retains $Z_{3}$ phase coherence, we call the intermediate phase “super”-fluid due to the phase coherence. Since $\cos(3\theta_{0})$ is relevant at the would-be BKT transition of a true $\tilde{N}$ superfluid [PhysRevB.16.1217], we expect that thermal transition of $\theta_{0}$ to be three-state Potts-like. Therefore, in either route, both transitions, between SN and the intermediate phase, and between the intermediate and the fully disordered phase, are in the three-state Potts universality class, reflecting $Z_{3}$ symmetry breaking at each step.
@@ VII Other fillings
Extending beyond $\nu=2/3$, the cluster constraint $\sum_{i\in\varhexagon_{r}}n_{i}\in\{0,1,2,3,4,5,6\}$ yields the “commensurate” fillings $\nu=0,1/3,2/3,1,4/3,5/3,2$. Since $\nu$ and $2-\nu$ are related by particle–hole symmetry and $\nu=0$ is trivial, it suffices to analyze $\nu=1/3$, $2/3$ and $1$. At $\nu=1/3$, Thm.  1 applies directly (no need for the tile set $\Sigma_{2}$), so $\tilde{N}$ remains an exact symmetry of the constrained Hilbert space. In contrast, at $\nu=1$ the leading ring exchange breaks $\tilde{N}$ (Fig.  9a). A constructive scheme (see SM Sec. C) can be proofed to generate all local operators order by order, yielding the leading ring exchanges (Fig.  9b) to be $\mathcal{O}_{1}$ for $\nu=1$, $\mathcal{O}_{2}$ for $\nu=2/3$, and $\mathcal{L}_{2}$ (the “Lemniscate operator” in Ref.[mao2023fractionalization]) for $\nu=1/3$. Notably, $\mathcal{L}_{2}$ conserves a sublattice dipole $\tilde{\vec{P}}=\sum_{i}n_{A}(x_{i})\vec{x}_{i}-\sum_{j}n_{A}(x_{j})\vec{x}_{j}$, which is violated by higher-order terms such as $\mathcal{L}^{\prime}$. Therefore we also expect the SN phase to exist at $\nu=1/3$. For $\nu=1$, the $\mathcal{O}_{1}$ term gaps the system into a sublattice charge-ordered state.
@@ VIII Summary and discussion
In summary, we uncovered a robust mechanism by which geometric frustration and local Hilbert-space constraints enforce an emergent, globally conserved quantum number for bosons under cluster charging interaction on the honeycomb lattice. By mapping the constrained Hilbert space at $\frac{2}{3}$ ($\frac{4}{3}$) or $\frac{1}{3}$ ($\frac{5}{3}$) filling to a nontrivial tiling problem and invoking the Conway–Lagarias boundary invariant, we rigorously established global conservation of the sublattice polarization $\tilde{N}=N_{A}-N_{B}$, beyond the previously shown $U(1)\times U(1)$ gauge symmetry of classical ground state[zhang2024bionicfractionalizationtrimermodel]. By mapping the local constraints to coupled double-dimers, we derive an effective field theory that faithfully incorporates the conservation of $\tilde{N}$ and the local constraints of the bosonic quantum Hamiltonian. Quantum fluctuations within this manifold give rise to an insulating yet macroscopically phase-coherent and gapless “supernematic” (SN) phase, characterized by a unique intertwined order that spontaneously breaks both the emergent global $U(1)$ symmetry and the lattice $C_{3}$ rotation, but not translation. We derived a non-local order parameter representing the SN that can be detected through ODLOR: the “bumped loop” operator, which is markedly different from familiar Wilson loop through its shape-dependent charge under $\tilde{N}$. We also analyzed the finite temperature transitions that sequentially restore the broken symmetries.
@@
We discover the SN phase from the field theory perspective based on expanding the cosine term around its minima. While this expansion is valid deep in the SN phase, its validity near the phase transition to the CO phase is an open question that requires further study. Nevertheless, the validity of the SN phase as a possible ground state can be further supported by considering a distinct microscopic route, without tuning quantum fluctuations starting from the CO phase. A concrete realization is provided by a trimer generalization of the Rokhsar–Kivelson Hamiltonian, which we refer to as the trimer–Rokhsar–Kivelson (trimer-RK) Hamiltonian. This model naturally extends the RK construction for quantum dimer models [RK] and includes both a ring-exchange term, $\mathcal{O}_{2}$, and a potential term (see Appendix E for explicit definitions). At the RK point, where the strengths of these two terms are equal, the ground state is an equal-weight superposition of all trimer-covering configurations, analogous to the dimer model at its RK point. The equal-time correlation functions at the trimer-RK point is the same as the ones in trimer liquid [zhang2024bionicfractionalizationtrimermodel]. In dimer models on bipartite lattices, such as the honeycomb lattice [moessner2001phase], perturbations away from the RK point toward stronger exchange terms drive the system into an ordered phase. In contrast, quantum fluctuations around the trimer-RK point generate an effective action that stabilizes the SN phase (see Appendix E). The key distinction between the proposed trimer-RK and dimer-RK Hamiltonians is rooted in the $\tilde{N}$ conservation. In the trimer case, the dynamics including stochastic fluctuations must preserve $\tilde{N}$ exactly, leading to a Langevin equation of the same form as that describing volume-conserving surface growth [Sun1989PhysRevA.40.6763]. Upon quantization, this yields the phase-angle Lagrangian identical to Eq. 23, in contrast to the quantum Lifshitz model that governs the dimer RK point on bipartite lattices [Henley2004J.Phys.:Condens.Matterc].
@@
As a physical platform to persue the realization of SN, we propose the Rydberg atom array. The Rydberg blockade mechanism has been shown to provide cluster-charging constraints on honeycomb lattices [kornjavca2023trimer]. The algebraic decay of various operators of the SN phase can be directly measured in this platform. Measuring the non-local order parameter can be challenging, but it may be achievable through strategies analogous to that used for topological phases in Ref. [Semeghini2021Sciencea]. Since the total number of particles is not fixed in Rydberg systems, we need to consider the stability of the SN under doping away from the commensurate fillings ($2/3$ or $1/3$). As was discussed in the thermal violation of $\tilde{N}$ conservation, the monomers induce a threefold anisotropy, which pins the macroscopic phase angle of the SN. Doping also introduces monomers that carry fractionalized charges of $e/3$ [mao2023fractionalization, zhang2024bionicfractionalizationtrimermodel]. However, since the $\vec{\eta}$ field is pinned in the SN phase, we expect these monomers to be confined at low temperature and to become deconfined only in the “super”-fluid regime at finite temperature.
@@
Acknowledgement. – We thank Jiaxin Qiao, Yizhi You, Cenke Xu, Yi-Zhuang You, Jing-Yuan Chen, Zheng-Cheng Gu, T. Senthil, Matthew Fisher, and Roderig Moessner for the stimulating discussions. We thank Christopher Mudry and Eduardo Fradkin for feedbacks on the draft. DM is supported by SNSF Grant No. 219339. E-AK was supported in part by the NSF through the grant OAC-2118310 supported by the U.S. Department of Energy through Award Number: DE-SC0023905. This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF10436 to E-AK and grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@ A.2 Symmetry properties under lattice transformations
In the continuum, the coarse-graining introduces non-zero average values of $h^{r,g,b}$, and the lattice symmetry can act with some additional shift on top of the microscopic translation [Fradkin2004Phys.Rev.Bb]. To see this clearly, we consider the nine colomnar ordering states (flat configurations with constant average colored height fields). The lattice transformation among these COs can be implemented via applying local operators across the lattice, allowing us to derive the values of height fields starting from one specific CO (see Fig.11). Hence, under lattice symmetries, the average heights transform as such,
@@ A.3 Relating coarse-grained fields to microscopic fields
However, this is not the full story due to the smearing in the coarse-graining. To the leading order in vertex operator, this amounts to match the lattice configuration of the colomnar ordering and the corresponding uniform heights in the continuum, analogous to dimer model [Fradkin2004Phys.Rev.Bb], which yields,
@@ Appendix B Sublattice density structure factor
Now let us consider the structure factor of the sublattice density $\tilde{n}(q)\equiv\sum_{i}n_{i,A}e^{iq\cdot x_{i,A}}-\sum_{j}n_{j,B}e^{iq\cdot x_{j,B}}$. From Eq.40, 42 and 45, we can write $\tilde{n}(q)$ in terms of $h_{0}$ fields. It is readily seen that
@@ Appendix E Perturbing around trimer-RK point
In the quantum dimer model on bipartite lattice, the dimer constraints can be viewed as Gauss law on the sites of the lattices [Fradkin2004Phys.Rev.Bb], and the low-energy physics can be described by compact $U(1)$ gauge theory in $2+1D$. The quantum fluctutation of the local moves in the dimer model corresponds to the instanton events and leads to confinement. Therefore, the gapless RK point is in general unstable under quantum fluctuation. We can ask the analogous question for the trimer model, that is, starting from the trimer RK point where the wavefunction is an equal weight superposition of all possible classical trimer configuration, what is the role of quantum fluctuation?
@@
The coefficient $K_{\eta}$ can be obtained by matching the Gaussian action of the classical trimer model in Ref. [zhang2024bionicfractionalizationtrimermodel]. In the dimer model near RK point, it was realized by Henley [Henley1997JStatPhysa, Henley2004J.Phys.:Condens.Matterc] that in the long wavelength limit, the plaquette flip dynamics can be captured by a Langevin equation of the height field (similar technique is also under the name of “stochastic quantization”, see [parisi1981perturbation, DIJKGRAAF2010365] and the reference therein). Let us only focus on the $h_{0}$ field. The dynamics of $h_{0}$ follows the following Langevin equation,
@@
which is of the same form as in volume conserving surface growth dynamics[Sun1989PhysRevA.40.6763].